OpenStudy (anonymous):

Select all of the following expressions that are equivalent to ab - 2 = c. a=C/B- 2 b=c+ 2/A AB=C+2 A-C/2=b B-2/C=a a=C+2/B

OpenStudy (anonymous):

Check for third option...

OpenStudy (anonymous):

b=c+ 2/A AB=C+2 a=C+2/B

OpenStudy (anonymous):